OpenStudy (anonymous):

@mathstudent55 May you help me with this question?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (imstuck):

I can help you, dear!

OpenStudy (aum):

Area of a triangle = 1/2 * base * height. Plug in the values and solve for x.

OpenStudy (aum):

Area = 400 square feet base = 40 ft. height = x Area of a triangle = 1/2 * base * height. 400 = 1/2 * 40 * x 400 = 20x x = ?

OpenStudy (valpey):

Seems like we need to know if their gardens are different sizes. If they are the same, you have your answer above.

OpenStudy (aum):

The houses are different sizes but the gardens are identical. Both have 40 ft. base and the two angles of one triangle are marked as equal to two angles of the other triangle.

OpenStudy (anonymous):

@aum Sorry for SUCH a late response, I was really tired yesterday. I got x = 20 . Would that be the final answer? ~Thank you for all who answered!

OpenStudy (anonymous):

@aum